OpenStudy (bekkah323):

simplify: 2/5+7/10 Times 3/2-4/15

OpenStudy (anonymous):

Numerator : 2/5 +7/10 = 4/10 +7/10 = 11/10 Denominator 3/2 -4/15 = 45/30-8/30=22/30 11/10-22/30=33/30-22/30=11/30

OpenStudy (anonymous):

o.0 Why the heck did he subtract it from each other when it said multiply =/ There's order of operations in here too. Is there parenthesis around 2/5 +7/10 and 3/2 - 4/15?

OpenStudy (anonymous):

@bekkah323